### v1

@@ 3 Discussion
In this work, an adaptive information-maximization encoding strategy is developed for ghost imaging (GI). The proposed framework is intended to serve as a generally applicable approach for information-optimal encoding in computational imaging systems with the encoding-decoding mode, provided that the forward imaging model can be explicitly characterized. In the present study, the forward model is formulated as a simplified linear model and demonstrated using a spatial-domain GI system with sequential bucket detection. In practical imaging scenarios, however, the forward process may be considerably more complex. Forward modeling of GI under realistic conditions, including effects such as optical blurring, down-sampling, and noise, has been extensively investigated in the literature [huang2020ghost, huang2021ghost, liu2025comprehensive]. Moreover, since Bayesian filtering theory is not restricted to linear systems, the proposed strategy can, in principle, be generalized to computational imaging problems governed by nonlinear forward models, such as phase retrieval and coherent detection. Beyond the form of the forward model, the framework is also not restricted to a specific sensing configuration. Although AIME is demonstrated with sequential bucket detection, the underlying Bayesian filtering and information optimization principles can be naturally extended to batch-measurement settings, which corresponds to computational imaging systems using two-dimensional sensor arrays. In such cases, multiple measurements are acquired simultaneously, and the framework can be generalized to adaptive or semi-adaptive encoding design in a higher-dimensional measurement space, suggesting its applicability to a broader class of computational imaging systems, like GI camera [liu2016spectral].
@@
Altogether, this work highlights the potential of information-theoretic encoding design as a principled approach to GI and, more broadly, to computational imaging systems. By explicitly modeling the information carried by detection signals under realistic physical constraints imposed by the imaging system, the proposed framework provides a systematic methodology to characterize and approach fundamental imaging limits. The finding that point-like encoding emerges as information-optimal under a total-energy constraint further offers a new information-theoretic perspective on the widespread use of point-to-point sampling strategies in biological vision systems. This result suggests that such sampling schemes may reflect an underlying principle of efficient information acquisition rather than merely engineering convenience. Beyond GI, the proposed framework is expected to inspire further information-theoretic studies on the evaluation and optimization of imaging and sensing systems, with potential implications for a broad range of applications, including autonomous sensing, machine vision, and data-efficient perception tasks. Importantly, in numerous such systems, flexible light-field modulation platforms such as metasurfaces and programmable photonic devices are increasingly employed as core encoding components, which imposes structured physical constraints on the encoding space. Incorporating those constraints into the proposed framework is supposed to offer a principled pathway that integrates emerging hardware platforms with the information-optimal encoding design.
@@
The optimization of the objective function is formulated as a constrained problem and solved using Projected Gradient Descent (PGD), which alternates between gradient descent and projection of parameters onto the constraint domain. Specifically, during the gradient descent step, we employ the SGD optimizer with learning rates of $1\times 10^{9}$ and $1\times 10^{5}$ for the mutual information and CRLB objective function, respectively. All other hyperparameters are set to the default values in PyTorch. To balance computational efficiency and estimation quality, we perform 100 PGD iterations per encoding optimization. All simulations are implemented in PyTorch 2.5.1 with CUDA 12.1 and executed on an NVIDIA GeForce RTX 3090 GPU. One epoch comprising both Bayesian filtering and encoding optimization takes approximately 0.2 s for a $128\times 128$ image.
@@
In the experiment, to determine the forward model shown as Eq. (1), the factor $\beta$ and the noise level $\{R_{k}\}$ are pre-measured. Specifically, a whiteboard with uniform reflectance is used as the imaging object for estimating $\beta$ and several Hadamard encoding patterns ${\bf h}^{(1)},\cdots,{\bf h}^{(M)}$ are pre-set. Then, the object is repetitively illuminated $L$ times with each pattern, and detection light intensity signals ${\bf z}^{(1)},\cdots{\bf z}^{(M)}$ are recorded, where each ${\bf z}^{(m)}\in{\mathcal{R}}^{L}$. And $\beta$ is estimated by
$\beta=\left\langle\frac{\overline{z}^{(m)}}{\sum_{i}{\bf h}^{(m)}_{i}}\right\rangle_{m}$
@@
where $\overline{z}^{(m)}$ denotes the average on elements of ${\bf z}^{(m)}$ and $\left\langle\cdot\right\rangle_{m}$ denotes the average on those estimation results of $\beta$ with $m=1,\cdots,M$ patterns.
For noise calibration of $R_{k}$, the detection noise variance, denoted as $\sigma^{2}_{(m)}$ corresponding to the $m$-th Hadamard pattern, is estimated using the sample variance over $L$ measurements:
$\sigma^{2}_{(m)}=\frac{\sum_{l=1}^{L}\left(z_{l}^{(m)}-\bar{z}^{(m)}\right)^{2}}{L-1},$ (14)
where $z_{l}^{(m)}$ represents the $l$-th measurement of the $m$-th pattern, and $\bar{z}^{(m)}$ is the sample mean. Consequently, the average signal-to-noise ratio (SNR) is defined as:
$\text{SNR}=10\log_{10}\left(\frac{\langle\bar{z}^{(m)}\rangle_{m}}{\langle\sigma^{2}_{(m)}\rangle_{m}}\right).$ (15)
The illumination intensity of the light source is adjusted to ensure the experimental SNR matches the target value. Given that the detection noise is modeled as $w^{2}{\bf H}{\bf x}$, the noise parameter $w^{2}$ for each pattern can be estimated via $w^{2(m)}=\sigma^{2}_{(m)}/\bar{z}^{(m)}$. The ensemble average is denoted as $\langle w^{2(m)}\rangle_{m}$.
In the numerical process, the acquired bucket detection values are normalized by a factor $\beta$. Therefore, the effective noise parameter used in the computation is scaled accordingly:
$w^{2}_{\text{eff}}=\frac{\langle w^{2(m)}\rangle_{m}}{\beta^{2}}.$ (16)
By varying the light intensity and repeating the aforementioned procedure, $w^{2}$ values corresponding to different detection intensity ranges are recorded. This parameter is essential for pattern optimization based on the Cramér-Rao Bound (CRB). In this work, we employ the same $w^{2}$ value as utilized in the mutual information-based optimization framework.
@@
Experimental setting in comparison with traditional imaging
In comparative experiments involving full-sampling conventional imaging under low SNR conditions, the dominant source of noise is set to arise from background light introduced with another LED source. Consequently, the background noise, denoted as $\sigma_{\text{back}}^{2}$, is modeled as additive Gaussian noise with constant variance. To quantify the SNR, $L$ measurements are acquired using a completely dark pattern. Let $z_{\text{back}}^{(l)}$ represent the signal intensity of the $l$-th measurement, and $\bar{z}_{\text{back}}$ denote the average intensity. The background noise variance $\sigma_{\text{back}}^{2}$ and the experimental SNR are estimated as follows:
@@
$\text{SNR}=10\log_{10}\left(\frac{\bar{z}_{\text{back}}}{\sigma_{\text{back}}}\right).$ (18)
@@
$p(a({\bf f}))=\frac{1}{2b({\bf f})}\exp\left[-\frac{\left|a({\bf f})-\mu({\bf f})\right|}{b({\bf f})}\right],$ (19)
where $\mu({\bf f})=0$ for those non-zero-frequency components, $b({\bf f})\propto(|{\bf f}|)^{-\gamma}$ with $\gamma\in[1,1.2]$. Based on Eq. (19), the parameters of the initial Gaussian PDF are set as
${\bf x}_{0}={\bf 0}$ (20a)
${\bf P}_{0}=\Psi^{\top}2{\bf b}{\bf b}^{\top}\Psi$ (20b)
where ${\bf b}=\left[b({\bf f}_{0}),b({\bf f}_{1}),\cdots\right]^{\top}$ and $\Psi$ denotes the digital cosine transform (DCT) matrix. It should be noted that, owing to the high dimensonality of image signals and the central limit theorem, although the frequency-domain coefficients are modeled as independent Laplacian variables, the image-domain pixels, being weighted superpositions of a large number of such coefficients, can be well approximated as following a Gaussian distribution.
Supplementary information
Materials include the Supplementary Text, Figs. S1 to S10, and Reference [1].
@@
C. Hu and S. Han conceived the idea and supervised the study. J. Sun and C. Hu performed most of the theoretical derivation. J. Sun and C. Hu conducted the experiment under discussion with Z. Bo, Z. Liu and M. Chen. L. Du and W. Liu provided the initial simulation codes for Bayesian filtering of GI, and contribute to code modification. J. Sun and C. Hu analyzed the experimental data, and wrote the original draft. All authors discussed the results contributed to editing and revising the paper.
@@ Appendix C Definition of the detection signal-to-noise ratio
Here we would like to discuss about the definition of the detection SNR in GI systems. Typically, the detection SNR in an imaging system is defined as
@@
where $\overline{z}$ represents the average value of the detection signal and $\sigma$ denotes the noise level. In the field of GI, there also exist some other kinds of definitions. For example, the bucket detection SNR (BSNR) is calculated as
${\rm BSNR}=10\log_{10}\left(\frac{\overline{z^{2}}-\overline{z}^{2}}{\sigma}\right),$ (35)
to measure the level of signal fluctuations. However, this definition only apply to GI systems with randomly fluctuating light fields, and it is not directly related to the detection process. In the field of signal processing, the SNR is defined as
where $P_{s}$ denotes the signal power and $P_{n}$ is the noise power. Different from traditional signal processing which deal with electronic signals, in GI the signal is in fact an optical one. Since the detector directly measures the intensity of light fields, the recorded intensity signal can already be considered as the power. Hence, in this study we still use the definition as Eq. (34).
@@
Figures A6 to A9 show experimental results of four representative imaging scenes for comparing GI with AIME-$\ell_{1}$ and fixed point-to-point conventional imaging.
(a) SNR about $6.86{\rm dB}$
(b) SNR about $8.47{\rm dB}$
Figure A6: Comparison results of scene 1 under different sampling ratios and SNRs, including imaging results and imaging quality
(a) SNR about $5.16{\rm dB}$
(b) SNR about $7.63{\rm dB}$
Figure A7: Comparison results of scene 2 under different sampling ratios and SNRs, including imaging results and imaging quality
(a) SNR about $5.90{\rm dB}$
(b) SNR about $7.63{\rm dB}$
Figure A8: Comparison results of scene 3 under different sampling ratios and SNRs, including imaging results and imaging quality
(a) SNR about $3.20{\rm dB}$
(b) SNR about $6.33{\rm dB}$
Figure A9: Comparison results of scene 4 under different sampling ratios and SNRs, including imaging results and imaging quality
@@ Appendix F Information-optimal encoding under different SNRs
Figures A14 and A15 show AIME results under three different SNRs, which indicates that AIME can adaptively deal with different imaging conditions and constraints.

### v2

@@ 3 Discussion
In this work, an adaptive information-maximization encoding strategy is developed for GI. The proposed framework is intended to serve as a generally applicable approach for information-optimal encoding in computational imaging systems with the encoding-decoding mode, provided that the forward imaging model can be explicitly characterized. In the present study, the forward model is formulated as a simplified linear model and demonstrated using a spatial-domain GI system with sequential bucket detection. While this simplified model enables analytical tractability and conceptual clarity, more realistic forward models accounting for optical blurring, down-sampling, and noise [huang2020ghost, huang2021ghost, liu2025comprehensive] can be further incorporated into the proposed framework. Moreover, since Bayesian filtering theory is not restricted to linear systems, the proposed strategy can, in principle, be generalized to computational imaging problems governed by nonlinear forward models, such as the coherent detection GI [pan2021micro]. Beyond the form of the forward model, the framework is also not restricted to a specific sensing configuration. Although AIME is demonstrated with sequential bucket detection, the underlying Bayesian filtering and information optimization principles can be naturally extended to batch-measurement settings. In such cases, multiple measurements are acquired simultaneously, and the framework can be generalized to adaptive or semi-adaptive encoding design in a higher-dimensional measurement space. This suggests its applicability to a broader class of computational imaging systems with array detection, like the GI camera [chu2021spectral].
@@
Altogether, this work highlights the potential of information-theoretic encoding design as a principled approach to GI and, more broadly, to computational imaging systems. By explicitly modeling the information carried by detection signals under realistic physical constraints imposed by the imaging system, the proposed framework provides a systematic methodology to characterize and approach fundamental imaging limits. The emergence of point-like encoding as information-optimal under a total-energy constraint offers an information-theoretic insight suggesting that point-to-point sampling strategies observed in biological vision [land2012animal] may reflect a principle of efficient information acquisition. Beyond GI, the proposed framework is expected to inspire further information-theoretic studies on the evaluation and optimization of imaging and sensing systems, with potential implications for a broad range of applications, including autonomous sensing, machine vision, and data-efficient perception tasks. Importantly, in numerous such systems, flexible light-field modulation platforms such as metasurfaces and programmable photonic devices are increasingly employed as core encoding components, which provides structured design space for encoding optimization. Incorporating this into the proposed framework is supposed to offer a principled pathway that integrates emerging hardware platforms with the information-optimal encoding design.
@@
The encoding optimization problem is formulated as a constrained optimization task and solved using Projected Gradient Descent (PGD), which alternates between gra- dient descent updates and projection onto the feasible constraint set. During the gradient descent step, we employ the SGD optimizer with learning rates of $1\times 10^{9}$ and $1\times 10^{5}$ for the mutual-information-based and CRB-based objectives, respectively. All remaining optimizer hyperparameters follow the default PyTorch configuration. To balance computational efficiency and estimation accuracy, each encoding optimization is performed using 100 PGD iterations. All numerical calculation are implemented in PyTorch 2.5.1 with CUDA 12.1 and executed on an NVIDIA GeForce RTX 3090 GPU. For a $128\times 128$ image, one complete cycle consisting of Bayesian filtering and encoding optimization requires approximately 0.12 s.
@@
The experiment is performed with a GI system with DMD modulation. To determine the forward model in Eq. (1), the scaling factor $\beta$ and the noise statistics $\{R_{k}\}$ are experimentally calibrated prior to imaging.
Calibration of $\beta$
@@
A whiteboard with uniform reflectance is used as the calibration object. A set of $M$ Hadamard encoding patterns ${\bf h}^{(1)},\cdots,{\bf h}^{(M)}$ is projected sequentially, and for each pattern $L$ repeated measurements are recorded, yielding detection vectors ${\bf z}^{(m)}\in{\mathcal{R}}^{L}$. Then, $\beta$ is estimated by
$\beta=\left\langle\frac{\overline{z}^{(m)}}{\sum_{i}{\bf h}^{(m)}_{i}}\right\rangle_{m}$ (14)
@@
where $\overline{z}^{(m)}$ denotes the average on elements of ${\bf z}^{(m)}$ and $\left\langle\cdot\right\rangle_{m}$ represents averaging over all calibration encoding patterns.
Calibration of noise statistics
As modeled in Section 2.1.1, the noise variance is parameterized as
$R_{k}=\omega^{2}{\bf h}_{k}^{\top}{\bf x}$
where $\omega^{2}$ characterizes the proportionality between signal intensity and noise variance and therefore controls the effective detection SNR. With those measurements on the whiteboard by using Hadamard patterns, estimation on $\omega^{2}$ in practical conditions is as follows. For the $m$-th encoding pattern, with $L$ repeated measurements, the corresponding noise coefficient is estimated as
${\omega^{2}}^{(m)}=\frac{{\sigma^{2}}^{(m)}}{{\bar{z}}^{(m)}}$ (15)
where ${\sigma^{2}}^{(m)}=\frac{\sum_{l=1}^{L}\left(z_{l}^{(m)}-\bar{z}^{(m)}\right)^{2}}{L}$ denotes the sample variance of the detection signal. The calibrated noise parameter is then taken as the ensemble average
$\omega^{2}_{\text{eff}}=\left\langle{\omega^{2}}^{(m)}\right\rangle_{m}.$ (16)
By adjusting the illumination intensity and repeating the calibration procedure, noise parameters corresponding to different operating SNR regimes are obtained. In each regime, the calibrated $\omega^{2}_{\text{eff}}$ is used consistently in both mutual-information-based and CRB-based encoding optimization.
@@
Experimental setting in comparison between GI with AIME-$\ell_{1}$ and conventional imaging
For experimental comparison between GI with AIME-$\ell_{1}$ and “fixed point-to-point” conventional imaging, both strategies are implemented using the same DMD-based GI system to ensure identical optical and detection conditions. For AIME-$\ell_{1}$, the optimized encoding patterns are physically realized with DMD. To improve illumination stability, each $4\times 4$ block of DMD micro-mirrors is grouped to form a single effective illumination pixel. The optimized patterns are mapped onto these grouped pixels and sequentially projected onto the scene, and the corresponding bucket detection signals are recorded. For conventional imaging, the same DMD-based GI system is operated in a raster-scanning mode. Specifically, a sequence of point-like illumination patterns is loaded onto the DMD, where each grouped $4\times 4$ block is illuminated one at a time following a fixed scanning order. The detected bucket signal associated with each illumination position is directly assigned as the grayscale value of the corresponding image pixel, thereby forming the image.
In the experiment under low-SNR condition, additional background noise is introduced using scattering light generated by a rotating ground glass. This produces a dynamic background whose temporal fluctuation, after bucket detection and time averaging, can be well approximated as additive Gaussian noise with constant variance. To quantify the noise level, $L$ measurements are acquired using a completely dark pattern. Let $z_{\text{back}}^{(l)}$ denote the $l$-th recorded signal, and $\bar{z}_{\text{back}}$ its sample mean over $L$ measurements, the background noise variance $\sigma_{\text{back}}^{2}$ is estimated as:
@@
With experimental detection signal $\{z_{k}\}$ corresponding to both AIME and raster-scanning, the experimental SNR is estimated by
$\text{SNR}=10\log_{10}\left(\frac{{\bar{z}}-\bar{z}_{\text{back}}}{\sigma_{\text{back}}}\right).$ (18)
where $\bar{z}$ is the average intensity of detection signals.
@@
$p(a({\bf f}))=\frac{1}{2\lambda({\bf f})}\exp\left[-\frac{\left|a({\bf f})-\mu({\bf f})\right|}{\lambda({\bf f})}\right],$ (19)
where $\mu({\bf f})$ for those non-zero-frequency components is set as 0, $\mu({\bf f}_{0})=\mu_{0}$ representing the zero-frequency amplitude is estimated with a uniform illumination, and $\lambda({\bf f})\propto(|{\bf f}|)^{-\gamma}$ with $\gamma\in[1,2]$. It should be noted that, owing to the high dimensonality of image signals and the central limit theorem, although the frequency-domain coefficients are modeled as independent Laplacian variables, the image-domain pixels, being weighted superpositions of a large number of such coefficients, can be well approximated to be subject to a Gaussian distribution. Based on this, the parameters of the initial Gaussian PDF are set as
${\bf x}_{0}=\mu_{0}{\bf 1}$ (20a)
${\bf P}_{0}=2\Psi^{\top}{\bf\Sigma}^{2}\Psi$ (20b)
where $\bf 1$ is an all-$1$ vector, $\bf\Sigma$} is a diagonal matrix whose diagonal elements consist of $\left\{\lambda({\bf f}_{0}),\lambda({\bf f}_{1}),\cdots\right\}$ and $\bf\Psi$ denotes DCT matrix.
@@
C. Hu and S. Han conceived the idea and supervised the study. J. Sun and C. Hu performed most of the theoretical derivation. J. Sun and C. Hu conducted the experiment under discussion with Z. Bo, Z. Liu and M. Chen. L. Du and W. Liu contribute to code modification for simulation and experimental data analyses. J. Sun and C. Hu analyzed the experimental data, and wrote the original draft. All authors discussed the results and contributed to editing and revising the paper.
@@ Appendix C Definition of the detection signal-to-noise ratio
Here we would like to discuss about the definition of the detection SNR in GI systems. In the field of GI, the bucket detection SNR (BSNR) [du2025information] is calculated as
${\rm BSNR}=10\log_{10}\left(\frac{\overline{z^{2}}-\overline{z}^{2}}{\sigma^{2}}\right),$ (34)
to measure the level of signal fluctuations. However, this definition only apply to GI systems with randomly fluctuating light fields, and it is not directly related to the detection process.
In the field of signal processing, the SNR is defined as
where $P_{s}$ denotes the signal power and $P_{n}$ is the noise power.
Different from traditional signal processing which deal with electronic signals, in GI the signal is in fact an optical one. Since the detector directly measures the intensity of light fields, the recorded intensity signal can already be considered as the power. Hence, in this study we use the definition as [bian2024broadband]
@@
where $\overline{z}$ represents the average value of the detection signal and $\sigma$ denotes the average noise level.
@@
Figures A6 to A9 show experimental results of four representative imaging scenes.
Figure A6: Comparison between GI with AIME-$\ell_{1}$ and conventional imaging on scene 1 under varying SRs and SNRs, including imaging results and imaging quality indicators. a1, a2, a3 are results under $4.85$ dB SNR, b1, b2, b3 are results under $7.03$ dB SNR.
Figure A7: Comparison between GI with AIME-$\ell_{1}$ and conventional imaging on scene 1 under varying SRs and SNRs, including imaging results and imaging quality indicators. a1, a2, a3 are results under $3.28$ dB SNR, b1, b2, b3 are results under $5.79$ dB SNR.
Figure A8: Comparison between GI with AIME-$\ell_{1}$ and conventional imaging on scene 1 under varying SRs and SNRs, including imaging results and imaging quality indicators. a1, a2, a3 are results under $3.89$ dB SNR, b1, b2, b3 are results under $5.79$ dB SNR.
Figure A9: Comparison between GI with AIME-$\ell_{1}$ and conventional imaging on scene 1 under varying SRs and SNRs, including imaging results and imaging quality indicators. a1, a2, a3 are results under $2.09$ dB SNR, b1, b2, b3 are results under $4.30$ dB SNR.
@@ Appendix F Information-optimal encoding under different SNRs
Figures A14 and A15 show AIME results corresponding to the same imaging scene pepper under three different SNRs, which indicates that AIME can adaptively deal with different noise levels and constraints, resulting in different encoding patterns.